OpenStudy (anonymous):

Write the equation, in standard form, of the line parallel to 4x− 7y = 7 that passes through (2, 9)

OpenStudy (texaschic101):

A parallel line will have the same slope. So we need to find the slope of 4x - 7y = 7. We will put this equation in y = mx + b form, where m is the slope. 4x - 7y = 7 -7y = -4x + 7 y = -4x/-7 + 7/-7 y = 4/7x - 1 the number in the m position, 4/7, is the slope. now we will use y = mx + b with slope(m) = 4/7 and points (2,9) y = mx + b 9 = 4/7(2) + b 9 = 8/7 + b 9 - 8/7 = b 63/7 - 8/7 = b 55/7 = b so your parallel equation is : y = 4/7x + 55/7 but we need it in standard form. Ax + By = C y = 4/7x + 55/7 -- multiply entire equation by 7 7y = 4x + 55 -- subtract 4x -4x + 7y = 55 -- multiply by -1 to make x positive 4x - 7y = -55 ===> standard form

OpenStudy (anonymous):

Thanks so much.

OpenStudy (texaschic101):

no problem :)